A body oscillates with simple harmonic motion according to the equation x(t)=5sin(4pie.t+pie/4

(x in meters). At time t = 5 seconds what are : /(6 x


2.5mrks)


(a) the displacement (b) the velocity

(c) the acceleration (d) the phase of the motion

(e) the frequency (f) the period of the motion.


1
Expert's answer
2021-10-06T10:04:39-0400

(a)


"x(5)=5\sin(4\pi(5)+\dfrac{\pi}{4})=2.5\sqrt{2}\ m"


(b)


"v(t)=20\pi\cos(4\pi t+\dfrac{\pi}{4})"

"v(5)=20\pi\cos(4\pi (5)+\dfrac{\pi}{4})=10\pi\sqrt{2}\ m/s"

(c)


"a(t)=-80\pi^2\sin(4\pi t+\dfrac{\pi}{4})"

"a(5)=-80\pi^2\sin(4\pi (5)+\dfrac{\pi}{4})=-40\pi^2\sqrt{2}\ m/s^2"

(d)


"\varphi(t)=4\pi t+\dfrac{\pi}{4}"

"\varphi(5)=4\pi (5)+\dfrac{\pi}{4}=\dfrac{81\pi}{4} rad"

(e)


"f=\dfrac{4\pi}{2\pi}=2\ s^{-1}"

(f)


"T=\dfrac{1}{f}=\dfrac{1}{2}\ s"